### v1

# Lead Times in Flux: Analyzing Airbnb Dynamics During Global Upheavals (2018-2022)

Abstract
Short-term changes in booking behaviors can significantly undermine naive forecasting methods in the travel and hospitality industry, especially during periods of global upheaval. Traditional metrics like average or median lead times capture only broad trends, often missing subtle yet impactful distributional shifts. In this study, we introduce a normalized L1 (Manhattan) distance to measure the full distributional divergence in Airbnb booking lead times from 2018 to 2022, with particular emphasis on the COVID-19 pandemic. Using data from four major U.S. cities, we find a two-phase pattern of disruption: a sharp initial change at the pandemic’s onset, followed by partial recovery but persistent divergences from pre-2018 norms. Our approach reveals shifts in travelers’ planning horizons that remain undetected by conventional summary statistics. These findings highlight the importance of examining the entire lead-time distribution when forecasting demand and setting pricing strategies. By capturing nuanced changes in booking behaviors, the normalized L1 metric enhances both demand forecasting and the broader strategic toolkit for tourism stakeholders, from revenue management and marketing to operational planning, amid continued market volatility.
Keywords: forecasting, lead time distribution, normalized L1 distance, COVID-19 pandemic, Airbnb, tourism analytics, consumer behavior, hospitality, revenue management. footnotetext:
1 Introduction
Booking lead time—the interval between the booking date and the actual stay date—is a key metric in the travel and hospitality industry. It influences operational and strategic decisions, including pricing strategies, inventory management, marketing campaigns, and customer relationship management. A thorough understanding of booking lead times enables businesses to optimize resource allocation, anticipate demand fluctuations, and enhance competitiveness in a dynamic market environment.
Prior research has underscored the importance of lead times in tourism management. Lim and McAleer (2002) showed how lead times affect revenue management strategies in the hospitality sector. Tussyadiah and Pesonen (2016) examined consumer decision-making processes in tourism, finding that booking lead times are influenced by a complex interplay of factors, including perceived risk, travel purpose, and technological advancements. Lead times are frequently incorporated into forecasts of aggregate demand (Fiori and Foroni, 2019; Pereira, 2016; Huang and Zheng, 2023; Zaki, 2022; Fiori and Foroni, 2020; Webb et al., 2022), while Katz et al. (2024) and de Oliveira et al. (2021) presented models for forecasting lead times directly.
Global events, particularly crises such as pandemics, have profound effects on travel behaviors and lead times. The COVID-19 pandemic, in particular, has led to unprecedented disruptions in the tourism industry (Sainaghi and Chica-Olmo, 2022; Sharma et al., 2021; Qiu et al., 2020; Mueller and Sobreira, 2024; Okafor et al., 2022). According to Gössling et al. (2021), the pandemic resulted in massive declines in international and domestic travel, altering consumer risk perceptions and booking behaviors. Sigala (2020) emphasized that COVID-19 not only impacted demand but also accelerated digital transformation and the adoption of flexible booking policies in the industry. Research suggests that travelers’ perceived risk and uncertainty lead to significant changes in booking patterns. For example, Lo et al. (2011) found that during health crises, travelers tend to delay bookings or opt for destinations perceived as safer. Fuchs and Reichel (2011) showed that risk reduction strategies, such as flexible cancellation policies and enhanced safety measures, become more critical in influencing booking lead times.
Despite recognition of these shifts, much of the existing literature relies on traditional statistical measures—such as mean and median lead times—to analyze booking behaviors. While these measures are useful, they may not fully capture the distributional changes and increased variability introduced by external shocks like pandemics. As Assaf and Tsionas (2019a, b, c) argued, there is a need for more advanced analytical approaches in tourism research to better understand complex consumer behaviors.
This study addresses this gap by introducing the normalized Manhattan (L1) distance metric as a novel approach to analyzing variations in booking lead time distributions. Unlike traditional metrics, the normalized L1 distance captures changes across the entire distribution, providing a more sensitive and comprehensive measure of divergence in booking behaviors over time. This method allows for the detection of shifts that might be overlooked when relying solely on averages or standard deviations.
The primary objective of this research is to assess the effectiveness of the normalized L1 distance metric in detecting and quantifying changes in Airbnb booking lead times during periods of global upheaval, specifically the COVID-19 pandemic. By analyzing data from Airbnb properties across four major U.S. urban centers—Austin, Boston, Miami, and San Francisco—from 2018 to 2022, we aim to provide a comprehensive perspective on the evolving dynamics of urban tourism.
Furthermore, this study contributes to the literature by offering empirical evidence on how external shocks impact booking behaviors in the sharing economy context. By situating our findings within the broader theoretical frameworks of risk perception and consumer behavior during crises, we enhance the understanding of how travelers adjust their planning in response to heightened uncertainty.
The normalized L1 distance is used to quantify the divergence in lead time distributions across different years, offering a non-parametric measure that adjusts for scale differences by normalizing the traditional L1 (Manhattan) distance. This approach gives a global summary of lead time distributional changes, allowing a more nuanced understanding of how booking patterns have shifted in response to significant events. This framework provides practical insights for industry stakeholders and also advances methodological approaches in tourism research.
2 Methodology
This section describes our approach to studying changes in Airbnb booking lead-time distributions from 2018 to 2022. In particular, we define the key variable of interest (§2.1), introduce the normalized L1 (Manhattan) distance for quantifying distributional divergences (§2.3), and outline our time-series decomposition and interpretative strategy (§2.4).
2.1 Data and Variable of Interest
We analyze Airbnb booking data for four major U.S. cities (Austin, Boston, Miami, and San Francisco) from January 2018 through December 2022. For each city, we consider both destination bookings (nights stayed in the city) and origin bookings (nights booked from the city), distinguishing between domestic and international travel corridors.
Monthly Lead-Time Distributions.
Define:
[…]
  • $\Delta\in\{0,1,\ldots,365\}$: the lead time in days between the booking date and the check-in date.
  • $\ell_{i}$: the number of nights in booking $i$.
  • $\Delta_{i}$: the lead time (in days) for booking $i$.
We first collect all bookings that start in month $t$. For each booking $i$, we note its lead time $\Delta_{i}$ and length of stay $\ell_{i}$. Then, for each $\Delta$,
$w_{t}(\Delta)\;=\;\sum_{\begin{subarray}{c}i:\\ \Delta_{i}=\Delta\end{subarray}}\ell_{i},$
is the total number of nights that began in month $t$ with a lead time of $\Delta$ days. [^1] In other words, if two guests each book 5-night stays with $\Delta_{i}=3$, that adds $5+5=10$ to $w_{t}(3)$. Since $\Delta$ ranges from 0 to 365, this yields $366$ lead-time categories. To express these counts as a probability distribution, we normalize:
$L_{t}(\Delta)\;=\;\frac{\,w_{t}(\Delta)\,}{\sum_{\delta=0}^{365}w_{t}(\delta)},\quad\Delta=0,1,\ldots,365.$ (1)
Thus, $L_{t}(\Delta)$ is the proportion of all nights that start in month $t$ which were booked $\Delta$ days in advance. By construction,
[…]
2.2 Descriptive Analysis
A descriptive analysis is performed for each city, examining each year separately. We look at:
  • Mean trip lead times: The mean trip lead time provides the average number of days between the booking date and the trip date, weighted by the trip length. It gives an overall sense of how far in advance travelers are planning their trips. The mean is calculated using the formula:
    $\text{Mean}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$
    where $n$ is the total number of trips and $x_{i}$ is the lead time for each trip.
  • Median trip lead times: The median trip lead time indicates the central tendency of trip lead times, offering a measure less sensitive to extreme outliers compared to the mean. This is particularly useful for understanding typical trip planning behavior in the presence of significant variability. It is calculated as follows:
    • If $n$ (the number of observations) is odd, the median is the middle value:
      $\text{Median}=x_{\left(\frac{n+1}{2}\right)}$
    • If $n$ is even, the median is the average of the two middle values:
      $\text{Median}=\frac{x_{\left(\frac{n}{2}\right)}+x_{\left(\frac{n}{2}+1\right)}}{2}$
      where $x_{(i)}$ represents the $i$-th ordered lead time.
  • Standard deviation of trip lead times: The standard deviation measures the variability in trip lead times, indicating how spread out the planning patterns are around the mean. This statistic helps to assess the consistency of trip planning behaviors over time. The standard deviation is computed as:
    Standard $\text{Standard Deviation}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\text{Mean})^{2}}$
    where $x_{i}$ is the lead time for each trip, and the mean is as defined above.
2.3 Divergence Analysis Using Normalized L1 Distance
While basic summary statistics (mean, median, and standard deviation) can highlight overall trends in lead times, they may overlook distributional shifts that arise during periods of upheaval. To address this, we employ the normalized L1 distance (a.k.a. the Manhattan distance).
Intuition:
The L1 distance between two probability distributions $p$ and $q$ on the same domain (in our case, $\Delta=0,1,\dots,365$) essentially sums the absolute differences of their probabilities at each point:
$D_{\text{L1}}(p,q)\;=\;\sum_{\Delta=0}^{365}\bigl{|}p(\Delta)-q(\Delta)\bigr{|}.$
If the distributions do not overlap at all (one is $1$ in places where the other is 00, and vice versa), the L1 distance is maximal. Because we want a measure in $[0,1]$, we divide by 2:
$D(p,q)\;=\;\frac{1}{2}\sum_{\Delta=0}^{365}\bigl{|}p(\Delta)-q(\Delta)\bigr{|}.$
Interpretation: $D(p,q)$ is the fraction of the total probability mass that has “shifted.” A value of 00 means no difference (identical distributions), and $1$ implies total divergence.
2.3.1 Year-over-Year vs. Baseline Comparisons
We calculate the normalized L1 distance between lead-time distributions in two ways:
  1. 1.
    Year-over-Year (YoY) Comparison:
    $l^{1}_{{\text{YoY}},t}\;=\;\frac{1}{2}\sum_{\Delta=0}^{\Delta_{\max}}\Bigl{|}L_{t}(\Delta)-L_{t-12}(\Delta)\Bigr{|}.$
    This measures how much the distribution in month $t$ diverges from the same calendar month in the previous year.
  2. 2.
    Baseline-2018 Comparison:
    $l^{1}_{{\text{YoY2018}},t}\;=\;\frac{1}{2}\sum_{\Delta=0}^{\Delta_{\max}}\Bigl{|}L_{t}(\Delta)-L_{2018,t}(\Delta)\Bigr{|},$
    where $L_{2018,t}(\Delta)$ denotes the distribution for the corresponding month of 2018. This approach captures how far month $t$ is from a pre-pandemic benchmark.
These two perspectives reveal distinct patterns. The YoY metric highlights sequential shifts (e.g., from 2020 to 2021), while the Baseline-2018 metric clarifies whether distributions are returning to pre-crisis norms or have settled into new equilibria.
2.4 Seasonal-Trend Decomposition (STL)
After computing the L1 distances over time, we apply a Seasonal-Trend decomposition using Loess (STL) (Cleveland et al., 1990) to each L1 time series. The STL method splits each series $y_{t}$ into three components:
$y_{t}=T_{t}+S_{t}+R_{t},$
where $T_{t}$ is a slowly varying trend, $S_{t}$ is a seasonal component capturing recurring monthly or weekly patterns, and $R_{t}$ is a remainder (irregular fluctuations). We prefer STL for its:
  • Robustness to Outliers: It uses iterative Loess smoothing to handle sudden spikes or dips.
  • Flexibility in Seasonal Patterns: Each city may show distinct seasonalities (e.g., academic schedules, major events).
  • Ease of Interpretation: We can pinpoint the long-term shift in lead-time divergences by examining $T_{t}$.
Using STL on $l^{1}_{{\text{YoY}},t}$ and $l^{1}_{{\text{YoY2018}},t}$ helps identify whether divergence spikes are driven mainly by structural (long-term) changes or by seasonal and event-driven fluctuations.
3 Toy Example: Impact of Lead Time Changes on Naive Forecast Methods
3.1 Setup
To illustrate the sensitivity of the normalized L1 metric to shifts in lead time distributions and to demonstrate how changes in lead times can affect commonly used forecast methods, we present a simulated example using a fictional town called B-Ville. In this example, we assume that in the year 2019, B-Ville had 1,000 total bookings for a specific trip date, and in 2020, it had 1,200 total bookings for the same trip date.
We created a lead time distribution—a compositional vector of size 30—representing the proportion of bookings made at each lead time from 1 to 30 days before the trip date. To simulate changes in booking behaviors from 2019 to 2020, we introduced a small perturbation to the 2019 lead time distribution by adding random noise drawn from a half-normal distribution with scale parameter $\sigma=0.05$ to each component and then re-normalizing the vector to ensure it sums to one. This resulted in the 2020 lead time distribution, reflecting a shift in booking patterns. The lead time distributions for both years are shown in Figure 1.
To estimate the pickup curves, we calculated the cumulative sum of the lead time distributions for each year. The pickup curves represent the accumulation of bookings as the trip date approaches, showing the cumulative percentage of total bookings made up to each lead time.
3.2 Results and Effects on the Pickup Method
The pickup method [^2] In this example, we apply the pickup method under the assumption of unconstrained demand. This means we consider that supply limitations do not affect the total number of bookings, allowing demand to be fully realized. is a common forecasting technique (Pereira, 2016) in the hospitality industry, where future bookings are predicted based on historical booking patterns and current observed bookings. For instance, suppose we are 17 days before the trip date in 2020 and have recorded 723 bookings so far. In 2019, the pickup curve indicates that by 17 days before the trip date, 69% of bookings had been made. Applying the multiplicative pickup method, we would predict the total number of bookings for 2020 to be
$\text{Forecasted Bookings}=\frac{\text{Current Bookings}}{\text{Cumulative \% at $t=17$ (from 2019 data)}}=\frac{723}{0.69}\approx 1,047.$ (2)
However, the actual total bookings in 2020 are 1,200. This means that the forecasted bookings underestimate the actual bookings by approximately 150 bookings, or 12.5%. This discrepancy arises because the lead time distribution shifted in 2020, with more bookings occurring closer to the trip date—a change captured by the normalized L1 distance between the 2019 and 2020 lead time distributions, which is 0.2156. This is illustrated in Figure 2, which displays both the historical pickup curves and the forecasted total bookings at each time point.
Traditional summary statistics fail to capture this change effectively. As shown in Table 1, the mean lead time increased by only 3.58%, the median remained unchanged, and the standard deviation decreased by 4.30%. These minor changes in conventional metrics mask the significant distributional shift affecting the pickup method’s forecast.
This example illustrates how changes in lead time distributions, as quantified by the normalized L1 metric, can impact the accuracy of forecasts made using historical lead times. When lead time distributions shift, relying solely on historical booking patterns may lead to underestimation or overestimation of future bookings.
3.3 Bounding the Forecast Error of Naive Methods Using the Normalized L1 Distance
We aim to derive a bound on the relative forecast error $\epsilon$ of the pickup method that incorporates both the normalized L1 distance $D$ between the historical and actual lead time distributions and the time remaining until the trip date.
Let:
  • $L_{\text{hist}}(k)$ and $L_{\text{actual}}(k)$ denote the historical and actual lead time distributions, where $k$ ranges from 1 to $\Delta_{\text{max}}$ (the maximum lead time considered).
  • $C_{\text{hist}}(\Delta)=\sum_{k=1}^{\Delta}L_{\text{hist}}(k)$ and $C_{\text{actual}}(\Delta)=\sum_{k=1}^{\Delta}L_{\text{actual}}(k)$ are the cumulative distributions up to lead time $\Delta$.
  • $D=\frac{1}{2}\sum_{k=1}^{\Delta_{\text{max}}}|L_{\text{hist}}(k)-L_{\text{actual}}(k)|$ is the normalized L1 distance between the distributions.
  • $\Delta_{\text{max}}$ is the maximum lead time considered (e.g., 365 days).
The relative forecast error $\epsilon$ of the pickup method is defined as:
$\epsilon=\frac{\hat{B}_{\text{total}}-B_{\text{total}}}{B_{\text{total}}}$ (3)
where:
  • $\hat{B}_{\text{total}}$ is the forecasted total number of bookings using the pickup method.
  • $B_{\text{total}}$ is the actual total number of bookings.
Using the pickup method, the forecasted total bookings are calculated as:
$\hat{B}_{\text{total}}=\frac{B_{\text{obs}}}{C_{\text{hist}}(\Delta)}$
where:
  • $B_{\text{obs}}$ is the number of bookings observed up to lead time $\Delta$.
  • $C_{\text{hist}}(\Delta)$ is the cumulative historical booking proportion up to lead time $\Delta$.
At time $\Delta$, the actual cumulative booking proportion is:
$C_{\text{actual}}(\Delta)=\frac{B_{\text{obs}}}{B_{\text{total}}}$
Substituting these expressions into Equation 3, we derive (see Appendix 6.1) the following bound on the relative forecast error $\epsilon$:
$|\epsilon|\leq\frac{2D\left(1-\frac{\Delta}{\Delta_{\text{max}}}\right)}{C_{\text{hist}}(\Delta)}$ (4)
This bound decreases as $\Delta$ increases (i.e., as we approach the trip date) and approaches zero when $\Delta=\Delta_{\text{max}}$.
3.3.1 Estimating the Normalized L1 Distance $D$
It is important to note that $D$ represents the divergence between the historical and actual lead time distributions, which is unknown in real time because the actual distribution $L_{\text{actual}}(k)$ is not fully observed until after the trip date. However, practitioners can estimate $D$ by:
  • Historical Analysis: Examining historical patterns of divergence during similar periods or events to estimate a typical or maximum value of $D$.
  • Scenario Analysis: Considering specific values of $D$ based on past maximum observed divergences to model potential forecast errors under different scenarios.
  • Real-Time Monitoring: Tracking early indicators of changes in booking behaviors (e.g., sudden shifts in booking pace) to adjust the estimated $D$.
By using these methods, forecasters can approximate $D$ and apply the bound in Equation 4 to assess the potential forecast error under various scenarios.
3.3.2 Example
Revisiting the toy example of B-Ville, suppose we estimate $D$ based on historical divergences:
  • Estimated Normalized L1 Distance: $D=0.2156$ (based on a previously observed maximum divergence).
  • Maximum Lead Time: $\Delta_{\text{max}}=30$ days.
  • At $\Delta=17$ days (13 days before the trip date):
    $1-\frac{\Delta}{\Delta_{\text{max}}}$ $=1-\frac{17}{30}=\frac{13}{30}$
    $C_{\text{hist}}(17)$ $=0.69$
    $|\epsilon|$ $\leq\frac{0.2156\times 2\times\frac{13}{30}}{0.69}\approx 0.27$
  • At $\Delta=25$ days (5 days before the trip date):
    $1-\frac{\Delta}{\Delta_{\text{max}}}$ $=1-\frac{25}{30}=\frac{5}{30}=\frac{1}{6}$
    $C_{\text{hist}}(25)$ $=0.90$
    $|\epsilon|$ $\leq\frac{0.215\times 2\times\frac{1}{6}}{0.90}\approx 0.08$
This example demonstrates how $D$ allows practitioners to use the bound to assess potential forecast errors at different lead times.
Note
The detailed derivation of this bound is provided in Appendix 6.1, where we explain how it arises from the properties of cumulative distributions and the normalized L1 distance.
4 Airbnb Results
4.1 Evolution of Lead Times
The analysis of Airbnb booking data from 2018 to 2022 shows significant variations in lead times across the cities of Austin, Boston, Miami, and San Francisco. We examined the mean, median, and standard deviation (SD) of lead times, segmented by domestic and international travel, and categorized by origin (bookings from the city) and destination (bookings to the city). Figures 3 and 4 show the trends in mean and median lead times, while tables 2 and 3 provide detailed yearly summaries.
4.1.1 2018–2019: Pre-Pandemic Period
During the pre-pandemic period, Miami consistently had the shortest median lead times for both inbound and outbound domestic travel, suggesting a higher propensity for last-minute bookings among travelers to and from Miami. For international travel, bookings originating from Miami and those destined for Austin had the shortest median lead times. In contrast, San Francisco had the longest median lead times for domestic and international destination bookings, indicating that travelers to San Francisco tended to plan their trips further in advance. On an origin basis, Boston had the longest median lead times for both domestic and international bookings, suggesting that residents of Boston planned their trips earlier than those from other cities.
4.1.2 2020: Impact of the COVID-19 Pandemic
The onset of the COVID-19 pandemic in 2020 caused significant disruptions in travel patterns, particularly in international travel. International destination median lead times surged by at least 20% in every city, with Boston and San Francisco experiencing the most substantial increases of 64% and 42%, respectively. The SD of lead times also reached its highest levels for international travel in 2020, reflecting heightened variability in booking behaviors.
Domestic travel had mixed responses. While Boston’s domestic destination median lead times increased by 17%, both San Francisco and Miami experienced decreases of 11% and 8%, respectively. This suggests that travelers to certain cities reacted differently to the pandemic’s onset. The SD for domestic travel lead times increased in several cities, indicating greater variability amid uncertainty.
4.1.3 2021: Recovery and Shifts in Booking Behavior
In 2021, as travel began to rebound, domestic lead times increased across all cities. Notably, Miami’s origin median lead time surged by 35%, and Austin’s destination median lead time increased by 21%. This suggests a return to more advanced planning for domestic trips. For international travel, lead times generally decreased in most cities, possibly reflecting pent-up demand and travelers’ eagerness to book trips as restrictions eased. San Francisco and Miami saw significant reductions in destination median lead times by 48% and 52%, respectively. Boston’s origin median lead time for international travel decreased by 29%.
4.1.4 2022: Stabilization and Continued Variability
By 2022, changes in mean and median lead times for domestic travel were generally less pronounced, suggesting a stabilization of booking behaviors. However, San Francisco experienced a notable 28% increase in destination median lead times, indicating that travelers to San Francisco were planning trips even further in advance than before. The SD reached its highest level for all origin domestic travel in 2022, implying persistent variability.
For international travel, mean and median lead times increased again for all cities, reflecting a cautious return to international travel with longer planning horizons. San Francisco experienced the largest year-over-year increases in origin mean and median lead times at 27% and 31%, respectively. Similarly, bookings to Miami saw median lead times increase by 23% and mean lead times by 22%.
4.2 Divergence Analysis Using Normalized L1 Distance
To quantify the changes in booking lead time distributions over time, we used the normalized L1 distance metric, calculating both the divergence from the baseline year 2018 ($l^{1}_{{\text{YoY2018}},t}$) and the year-over-year divergence ($l^{1}_{{\text{YoY}},t}$). Figures 5-8 display these divergences, with detailed statistics summarized in Tables 8-11.
4.2.1 Baseline Comparison Divergence ($l^{1}_{{\text{YoY2018}},t}$)
The divergence from the baseline year 2018 shows significant shifts in booking behaviors due to the pandemic. For international travel, divergence peaked in 2020 for all cities, indicating substantial disruptions. For example, the median $l^{1}_{{\text{YoY2018}},t}$ value for international bookings to Boston increased from 0.05 in 2019 to 0.34 in 2020. Similarly, for bookings originating from Austin, the median divergence increased from 0.05 in 2019 to 0.17 in 2020. Although there was some reduction in divergence in 2021 and 2022, the values remained higher than pre-pandemic levels, suggesting continued adjustments in international travel behaviors.
For domestic travel, divergence from the 2018 baseline increased in 2020 and remained elevated through 2022. For instance, the median $l^{1}_{{\text{YoY2018}},t}$ for domestic bookings from San Francisco increased from 0.03 in 2019 to 0.11 in 2020 and remained around 0.11–0.12 in subsequent years. In Boston, the divergence for domestic destination bookings increased from 0.03 in 2019 to 0.10 in 2020, rising further to 0.15 in 2022. These sustained divergences indicate prolonged changes in domestic booking behaviors compared to the pre-pandemic period.
4.2.2 Year-over-Year Divergence ($l^{1}_{{\text{YoY}},t}$)
The year-over-year divergence analysis shows significant temporal shifts in booking patterns. For domestic travel, $l^{1}_{{\text{YoY}},t}$ values peaked in 2021 for all cities, reflecting ongoing adjustments even as travel began to recover. For example, the median $l^{1}_{{\text{YoY}},t}$ for domestic bookings to San Francisco increased from 0.04 in 2019 to 0.12 in 2020 and further to 0.19 in 2021. In Austin, the median $l^{1}_{{\text{YoY}},t}$ for domestic bookings from the city increased from 0.04 in 2019 to 0.09 in 2020 and 0.13 in 2021.
For international travel, divergences peaked in either 2020 or 2021. The median $l^{1}_{{\text{YoY}},t}$ for international bookings from Miami increased from 0.04 in 2019 to 0.13 in 2020, with a slight decrease to 0.11 in 2021. For bookings to Boston, the median divergence increased from 0.05 in 2019 to 0.33 in 2020 and 0.39 in 2021. By 2022, divergences began to decrease across all cities, indicating a gradual stabilization of booking behaviors.
These patterns underscore the disruptions caused by the COVID-19 pandemic and the varied recovery trajectories across different cities and travel types. The elevated divergences in 2021 compared to 2020 for domestic travel show a ”two-phase” anomaly phenomenon, where the anomalous year 2020 serves as a new baseline, leading to higher $l^{1}_{{\text{YoY}},t}$ values even as travel patterns start to stabilize.
4.2.3 Correlation Analysis of Lead Time Divergences
The correlations between domestic and international $l^{1}$ divergences across cities and over time are shown in figures 9 and 10. High correlations within the same travel type and direction suggest consistent behavioral shifts, while lower correlations between different travel types indicate distinct factors influencing domestic and international booking patterns.
In Miami, there was a strong correlation between domestic and international divergences, suggesting that local conditions similarly affected both travel segments. For instance, Miami’s domestic origin $l^{1}_{{\text{YoY2018}},t}$ had a correlation of 0.83 with its international destination divergence. In contrast, San Francisco exhibited lower correlations between domestic and international travel, implying that different factors influenced booking behaviors in these segments.
Lag effects were also observed, particularly in Austin, where domestic trends influenced international trends with a delay (correlation of 0.76 in $l^{1}_{{\text{YoY2018}},t}$). Consistently high correlations within categories across cities indicate a degree of interconnectedness in travel patterns, with domestic travel patterns significantly influencing international travel in some cases.
4.2.4 Seasonal-Trend Decomposition (STL) Analysis
The Seasonal-Trend decomposition using LOESS (STL) on the normalized L1 distance time series for both the baseline year 2018 ($l^{1}_{{\text{YoY2018}},t}$) and the year-over-year ($l^{1}_{{\text{YoY}},t}$) divergences are shown in figures 1114. The trend components from the STL give insights into the persistent changes in booking lead time distributions over time.
For the $l^{1}_{{\text{YoY2018}},t}$ divergences, the trend components for both domestic and international travel peaked around mid-2020 to mid-2021, corresponding with the peak impact of the COVID-19 pandemic on global travel. These trend components do not return to zero by 2022, indicating that booking lead time distributions have not reverted to their pre-pandemic (2018) patterns. For example, the trend components for domestic origin divergences in Boston and San Francisco remain elevated going into 2023, suggesting sustained shifts in booking behaviors in these markets. This persistent divergence implies that travelers continue to book accommodations with lead times significantly different from those in 2018, reflecting lasting changes in planning behavior possibly due to altered risk perceptions, remote work arrangements, or new travel preferences.
In contrast, the $l^{1}_{{\text{YoY}},t}$ divergences show a different pattern. For domestic travel, the trend components generally decrease towards pre-pandemic levels by 2022, indicating that year-over-year changes in booking behaviors are stabilizing. However, the divergences remain higher than zero, suggesting that while the rate of change is normalizing, booking patterns have not fully returned to previous norms. For international origin travel, the trend components show an acceleration of divergence going into 2023. This indicates that international booking behaviors continue to change significantly compared to the immediate past year.
These findings show that booking lead time distributions have undergone significant and lasting changes since the onset of the pandemic. These changes are not uniformly reverting to pre-pandemic patterns, especially in international travel, showing the importance of using comprehensive divergence metrics to inform industry responses and strategies.
5 Discussion
Our analysis showed differential recovery patterns among the cities studied, contrasting the influence of local policies and destination-specific factors on tourism resilience and booking behaviors. Cities like Miami and Austin exhibited quicker recoveries in domestic travel lead times, returning closer to pre-pandemic patterns by 2022. This resilience may be attributed to less stringent local COVID-19 restrictions, a higher proportion of leisure travelers, and their appeal as destinations—factors that align with prior research indicating that destinations with flexible policies and strong domestic appeal recovered more swiftly from pandemic-induced downturns (Sigala, 2020). In contrast, Boston and San Francisco showed prolonged divergences in both domestic and international travel lead times, likely due to stricter local regulations, reliance on international business travel, and cautious reopening strategies. These observations are consistent with studies noting that cities dependent on international and business travel faced greater challenges during the pandemic (Götz et al., 2021).
These findings align with theories of risk perception and consumer behavior during crises, suggesting that uncertainty and perceived risk lead to more cautious consumer behaviors, such as delayed booking and a preference for flexible options (Lo et al., 2011). The increased lead times and heightened variability during the pandemic years corroborate this, particularly in international travel, which experienced greater disruptions due to heightened risk perceptions and fluctuating international policies. The impact of global crises on travel patterns has been extensively researched, and our study extends this understanding by showing how these factors influence not only overall visitation numbers but also the temporal dynamics of booking behaviors and forecasting accuracy.
Furthermore, the limitations of traditional summary statistics in capturing these behavioral shifts underscore the need for advanced analytical approaches in tourism research. By introducing the normalized L1 distance metric, our study contributes to methodological advancements in understanding complex consumer behaviors. This metric effectively quantified changes in booking lead times and showed shifts not apparent through traditional measures, enhancing our ability to understand and forecast traveler behavior in response to external shocks. As our toy example demonstrated, even small divergences captured by the normalized L1 distance can lead to significant under- or overestimation of demand. The derived forecast error bound offers a practical tool to assess and mitigate these errors by estimating the normalized L1 distance and adjusting forecasting models accordingly.
The sustained divergences in booking behaviors have important implications for the tourism and hospitality industry. Businesses should anticipate continued changes in lead time distributions and not expect a return to pre-pandemic patterns. These shifts suggest that traditional forecasting models based on historical data may no longer suffice, necessitating updates to demand forecasting and revenue management strategies. Practitioners relying on naive forecasting methods, such as the pickup method, should be aware of potential forecast errors due to shifts in lead time distributions. Moreover, the increasing trend in year-over-year divergences for international origin travel implies that booking behaviors are still evolving, underscoring the importance of flexibility and adaptability in business operations. Companies should monitor these trends closely to respond promptly to changes in traveler behavior, such as offering flexible booking policies or tailoring marketing efforts to match current planning horizons.
While our study offers valuable insights, it has limitations. The analysis is confined to four U.S. cities and one accommodation platform, which may limit generalizability. Future research could expand to include a broader range of destinations and accommodation types. Additionally, ongoing developments, such as changes in remote work policies and evolving traveler preferences, may continue to impact booking behaviors. Longitudinal studies extending beyond 2022 would help in understanding the long-term effects of the pandemic and other global events on tourism. Further methodological advancements could enhance the applicability of the normalized L1 distance metric and the derived forecast error bounds. Integrating the normalized L1 distance into predictive models or combining it with other metrics may offer more robust tools for industry stakeholders.
Acknowledgement
The authors thank Sean Wilson, Kai Brusch, Liz Medina, and Jenny Cheng for helpful discussions, Ellie Mertz for championing the research, and Lauren Mackevich and Lori Callo for their indispensable operational support.